# Regression-Based Proximal Causal Inference

### v1

@@ A.6.3 Log $Y$ Link and Logit $W$ Link Combination with No Interaction Term
$U=\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(W=0);$
$*.\displaystyle|E[U|A,Z,W=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.}$.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,W=0]=t(A,Z)$, i.e.,
@@
A.6.4 Log $Y$ Link and Logit $W$ Link Combination with Interaction Terms
@@
$U=\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|(W=0);$
$*.\displaystyle|E[U|A,Z,W=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.}$.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,W=0]=t(A,Z)$, i.e.,
@@
A.7 Simulation Study
@@
$*.\displaystyle|E[U|A,Z|Y=0,W=0]|<\infty,\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0,W=0$ exists for any real value *.}$.
@@
This result allows for generating $(Y,W)$ from $(U,A,Z)$ through $f(Y,W|U,A,Z)$ via the accept-reject sampling algorithm.
@@
    Generating $(Y,W)$ from $f(Y,W|U,A,Z)$ presented in (97) with $\beta_{0}=-1.4$, $\beta_{a}=1.2$, $\beta_{u}=-0.7$, $\beta_{w}=\alpha_{y}=0.5$, $\alpha_{0}=-0.8$ and $\alpha_{u}=0.5$ via the accept-reject sampling algorithm.
@@
A.8 Statistical Properties of the Two-stage Estimators
@@
where $g_{1}$ and $g_{2}$ are link functions chosen according to the nature of $W$ and $Y$, and $S=S(\alpha^{*})$ stresses $S$ is a function of parameters $\alpha^{*}$. The asterisk superscript (***) is defined in Table 3. The displayed regression models only involved main effects, but the derivation follows similarly with models involving interaction or transformed variables.
Table 1 lists the regression models for cases involving logit links. Although the notation below does not apply to cases where the logit link is involved, the inference of the parameters can be derived following the same approach.

### v3

@@ A.7.3 Log $Y$ Link and Logit $W$ Link Combination with No Interaction Term
$U|A,Z,W=0\sim\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|W=0;$
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,W=0]=t(A,Z)$, i.e.
@@
A.7.4 Log $Y$ Link and Logit $W$ Link Combination with Interaction Terms
@@
$U|A,Z,W=0\sim\underbrace{E[U|A,Z,W=0]}_{=h(A,Z)}+\epsilon;\ E[\epsilon]=0;\ \epsilon\perp\!\!\!\perp(A,Z)|W=0;$
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|W=0$ exists for any real value *.
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z,W=0]=t(A,Z)$, i.e.
@@
By similar algebra steps, we can solve the case where the logit link is applied to the multinomial $Y$ or $W$. The procedures are given in the two gray cells of the table below. Details are omitted.
A.8 Simulation Study
@@
This result allows for generating $(Y,W)$ from $(U,A,Z)$ through $f(Y,W|U,A,Z)$ via a multinomial distribution (which encodes the bivariate Bernoulli distribution using 4 levels).
@@
    Generating $(Y,W)$ from $f(Y,W|U,A,Z)$ presented in (99) with $\beta_{0}=-1.4$, $\beta_{a}=1.2$, $\beta_{u}=-0.7$, $\beta_{w}=\alpha_{y}=0.5$, $\alpha_{0}=-0.8$ and $\alpha_{u}=0.5$ via a multinomial distribution which encodes the corresponding bivariate Bernouli distribution using 4 levels.
@@
The codes for simulation in R are available at: https://github.com/NoBirdInTree/Regression-Based-PCI/tree/main.
@@
A.9 Statistical Properties of the Two-stage Estimators
@@
where $g_{1}$ and $g_{2}$ are link functions chosen according to the nature of $W$ and $Y$, and $S=S(\alpha^{*})$ stresses $S$ is a function of parameters $\alpha^{*}$. The asterisk superscript (*) is defined in Table 2. The displayed regression models only involved main effects, but the derivation follows similarly with models involving interaction or transformed variables.
Table LABEL:table:1 lists the regression models for cases involving logit links. Although the notation below does not apply to cases where the logit link is involved, the inference of the parameters can be derived following the same approach.
@@
A.10 R & SAS Codes Demonstration for the 2-Stage GLM
In this section, we provide annotated R and SAS codes for the two-stage GLM Procedures 1, 2, and 3 established in the main text. Note that this is only for estimation purposes; for variance estimation, bootstrap or M-estimation methods should be used. For R implementation, please refer to the pci2s package, for which we attach a brief demonstration at the end.
R Codes:
# Load the data into a dataframe
mydata <- data.frame(
  y = c(1, 0, 1, 0, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0),
  a = c(0, 0, 1, 0, 1, 1, 1, 0 ,1, 0, 1, 1, 1, 1, 0),
  z = c(3, 4, 5, 6, 1, 2, 4, 1, 1, 4, 3, 7, 1, 3, 3),
  w = c(1, 0, 0, 0, 1, 1, 1, 0 ,0, 1 ,0, 1, 0, 1, 1)
# Linear (Y, W) Case
# Perform the first-stage linear regression W ~ A + Z
m1 <- lm(w ~ a + z, data = mydata)
# Compute the estimated E[W|A,Z]
S <- m1\$fitted.values
# Perform the second-stage linear regression Y ~ A + E[W|A,Z]
m2 <- lm(y ~ a + S, data = mydata)
summary(m2)
# Count (Y, W) Case
# Perform the first-stage Poisson regression W ~ A + Z
m1 <- glm(w ~ a + z, family = poisson(), data = mydata)
# Compute the estimated log(E[W|A,Z])
S <- m1\$fitted.values
# Perform the second-stage Poisson regression Y ~ A + log(E[W|A,Z])
m2 <- glm(y ~ a + S, family = poisson(), data = mydata)
summary(m2)
# Binary (Y, W) Case
# Perform the first-stage logistic regression W ~ A + Z + Y
m1 <- glm(w ~ a + z + y, family = binomial(), data = mydata)
# Compute the estimated logit(P(W=1|A,Z,Y=1))
S <- m1\$coefficients[1] + m1\$coefficients[2] * mydata\$a +
  m1\$coefficients[3] * mydata\$z + m1\$coefficients[4]
# Perform the second-stage logistic regression Y ~ A + logit(P(W=1|A,Z,Y=1)) + W
m2 <- glm(y ~ a + S + w, family = binomial(), data = mydata)
summary(m2)
SAS Codes:
data mydata;
    input y a z w;
    datalines;
    1 0 3 1
    0 0 4 0
    1 1 5 0
    0 0 6 0
    1 1 1 1
    1 1 2 1
    0 1 4 1
    1 0 1 0
    1 1 1 0
    1 0 4 1
    1 1 3 0
    1 1 7 1
    1 1 1 0
    1 1 3 1
    0 0 3 1
run;
* Linear (Y, W) Case;
* Perform the first-stage linear regression W ~ A + Z;
proc reg data=mydata;
    model w = a z;
    output out=first_stage p=S;
run;
* Perform the second-stage linear regression Y ~ A + E[W|A,Z];
proc reg data=first_stage;
    model y = a S;
run;
* Count (Y, W) Case;
* Perform the first-stage Poisson regression W ~ A + Z;
proc genmod data=mydata;
    model w = a z / dist=poisson link=log;
    output out=first_stage p=S;
run;
* Perform the second-stage Poisson regression Y ~ A + log(E[W|A,Z]);
proc genmod data=first_stage;
    model y = a S / dist=poisson link=log;
run;
* Perform the first-stage logistic regression W ~ A + Z + Y;
proc logistic data=mydata;
    model w(event=’1’) = a z y;
    ods output ParameterEstimates=logistic_model;
run;
* Extract the coefficients from the logistic model;
proc sql noprint;
    select estimate into :gamma_it from logistic_model where variable=’Intercept’;
    select estimate into :gamma_a from logistic_model where variable=’a’;
    select estimate into :gamma_z from logistic_model where variable=’z’;
    select estimate into :gamma_y from logistic_model where variable=’y’;
quit;
* Compute the estimated logit(P(W=1|A,Z,Y=1));
data new_data;
    set mydata;
    S = &gamma_it + a * &gamma_a + z * &gamma_z + &gamma_y;
run;
* Perform the second-stage logistic regression Y ~ A + logit(P(W=1|A,Z,Y=1)) + W;
proc logistic data=new_data;
    model y(event=’1’) = a S w;
run;
R demonstration of for using pci2s package:
# ?devtools::install_github
devtools::install_github("KenLi93/pci2s")
library(pci2s)
# Load the data into a dataframe
Y <- c(1, 0, 1, 0, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0)
A <- c(0, 0, 1, 0, 1, 1, 1, 0 ,1, 0, 1, 1, 1, 1, 0)
Z <- c(3, 4, 5, 6, 1, 2, 4, 1, 1, 4, 3, 7, 1, 3, 3)
W <- c(1, 0, 0, 0, 1, 1, 1, 0 ,0, 1 ,0, 1, 0, 1, 1)
# Linear (Y, W) Case
p2sls_lm_results <- p2sls.lm(Y = Y, A = A, Z = Z, W = W, Xw = cbind(1, A, Z))
p2sls_lm_results\$summary_first_stage
p2sls_lm_results\$summary_second_stage
# Count (Y, W) Case
p2sls_loglin_results <- p2sls.loglin(Y = Y, A = A, Z = Z, W = W, Xw = cbind(1, A, Z))
p2sls_loglin_results\$summary_first_stage
p2sls_loglin_results\$summary_second_stage
# Binary (Y, W) Case
p2sls_logit_reg_results <- p2sls.logitreg(Y = Y, A = A, Z = Z, W = W)
p2sls_logit_reg_results\$summary_first_stage
p2sls_logit_reg_results\$summary_second_stage